# Mechanistic Modeling of Continuous Lyophilization for Pharmaceutical Manufacturing

### v1

@@ Abstract
Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products, e.g., mRNA vaccines, in pharmaceutical manufacturing. While extensive efforts have been dedicated to shift the pharmaceutical industry towards continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in a batch mode. This article proposes the first mechanistic model for a complete continuous lyophilization process, which includes freezing, primary drying, and secondary drying. The state-of-the-art lyophilization technology is considered, in which vials are suspended and moved continuously through the process. The model can describe the evolution of several critical process parameters, namely the product temperature, ice/water fraction, sublimation front position, and concentration of bound water, for the entire lyophilization process. The model is also demonstrated for several applications related to process design and optimization. Ultimately, the framework and results presented in this work can serve as a solid foundation to guide the design and development of future continuous lyophilization processes.
@@
Keywords— Lyophilization, Freeze Drying, Continuous Manufacturing, Continuous Lyophilization, Pharmaceutical Manufacturing, Suspended Vials
@@ Introduction
Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products in pharmaceutical manufacturing [1]. Various promising applications of lyophilization have been proven, including its recent applications that provide long-term stability for mRNA vaccines (e.g., for COVID-19) to allow for higher storage temperature [2, 3]. Advancement in lyophilization technology could therefore play a crucial role in the future of the pharmaceutical industry. A typical lyophilization process consists of three main steps, namely (1) freezing, (2) primary drying, and (3) secondary drying. During freezing, the product and liquid (usually water) in a vial are cooled such that most of the liquid (free water) is frozen, with a small amount of liquid (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals. In primary drying, the free water in the form of ice crystals is removed via sublimation under vacuum. Finally, in secondary drying, the product is heated further such that the bound water can be removed via desorption.
@@
Although the current trends in the pharmaceutical industry largely focus on the adoption of continuous manufacturing, the majority of production-scale lyophilization is still being operated in a batch mode [4, 5], with most research efforts dedicated to process optimization, monitoring, and control to ensure that the final product quality meets the regulations [6, 5]. Conventional lyophilization of unit doses usually entails a batch of vials containing drug products situated on the cooling/heating shelf where the entire lyophilization process occurs. A number of continuous lyophilization concepts for pharmaceutical products have been proposed, including spray freeze drying [7] and spin freezing [4] (see a comprehensive review of continuous lyophilization technologies in Ref. [5]). A novel lyophilization technology has been recently proposed by Ref. [6], in which vials are suspended and continuously moved along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
@@
Mathematical models have been widely used to assist the design, optimization, and control of lyophilization processes. Since the phenomena in lyophilization are underpinned by several theories in the area of heat and mass transfer, models for lyophilization are mostly physics-based rather than data-driven. Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [8, 9, 10, 11, 12], primary drying [13, 14, 15, 16, 17, 8, 18, 19, 20, 21, 22], and secondary drying [13, 14, 15, 16, 17, 23, 21, 24, 25]. Only a few models for continuous lyophilization are available. For example, a model for the primary drying step in continuous drying of spin frozen vials was proposed by Ref. [26]. A model for the freezing step was also developed for spin freezing [27]. For spray freeze drying, Ref. [28] developed a detailed model that can predict the temperature of droplets during the cooling and freezing phases. Nevertheless, there is no model for the suspended-vial configuration. Besides, none of the published models for continuous lyophilization considers a complete lyophilization cycle (all three steps), which is critical for optimization and control of the entire continuous operation.
@@
This article presents the first mechanistic model for continuous lyophilization of suspended vials. The model is developed to capture important phenomena of the entire lyophilization process, including cooling and stochastic/controlled ice nucleation during the freezing step, heat and mass transfer with sublimation during the primary drying step, and heat transfer with desorption during the secondary drying step. The proposed model is validated and employed to predict the evolution of important process parameters, with several applications associated with process design and optimization demonstrated.
@@
This article is organized as follows. Section 2 gives an overview of conventional (batch) and continuous lyophilization technologies. Section 3 extensively describes the development of our mechanistic model and its implementation, supported by a comprehensive review on a variety of modeling strategies used in the literature. Section 4 validates the proposed model with the experimental data from the relevant literature. Section 5 showcases four case studies to demonstrate various applications of our model. Finally, Section 6 summarizes the study and briefly discusses possible future work.
@@
The majority of lyophilization processes in the pharmaceutical industry has been operated in a batch mode for many decades. The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Fig. 1A). The bottom shelf, whose temperature can be manipulated, is used to cool the vials during the freezing step and heat the vials during the drying steps.
Figure 1: (A) Conventional batch lyophilization of unit doses. A number of vials are placed on the cooling/heating shelf. (B) Continuous lyophilization of suspended vials. A number of vials are suspended and moved continuously along the lyophilizer.
@@
A comprehensive review of continuous lyophilization technologies is given in Ref. [5]. To summarize, the current continuous lyophilization technologies can be divided into four main categories. The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [29] or a revolving plate with holes [30]. Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with high surface-to-volume ratio, greatly improving heat and mass transfer in the process [7, 28]. Both aforementioned approaches are proposed for lyophilization of bulk material. For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [4].
Finally, the current state-of-the-art continuous lyophilization technology for unit doses has been recently proposed by Ref. [6]. This technology employs the suspended-vial configuration as illustrated in Fig. 1B, where all vials are suspended and continuously moved through the freezing and drying chambers, without any contact between the vials and shelf. This suspended-vial technology has several benefits over the existing ones. First, heat transfer in the equipment is uniform because every vial experiences the exact same heat transfer conditions along the way. In addition, there is a dedicated chamber for controlling the ice nucleation process to help reduce variation in the crystal structure caused by the stochastic nature of ice nucleation. Second, the process does not involve with complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control. Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in pharmaceutical applications. Finally, with the automatic filling and load-lock systems, the process is fully continuous. Therefore, our mechanistic model developed in this work is based on this suspended-vial configuration. We refer to Ref. [6] for detailed information and discussion related to the equipment design.
@@
This section develops a mechanistic model for continuous lyophilization based on the suspended-vial configuration (see Fig. 1B for a simple schematic and Ref. [6] for detailed discussion). The section starts by discussing the existing models in the literature and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
Review of the existing models
Mechanistic modeling for conventional batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry. Despite the fact that the models are for batch processes, modeling strategies underpinning those models can be used as a basis for the development of a continuous lyophilization model for suspended vials. Although there are some previous studies briefly discussing the literature on mechanistic modeling of lyophilization [18, 21], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation. Hence, this section firstly gives a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes the optimal model strategies to be used. Note that published models for continuous lyophilization mentioned in Section 1 are based on the spin freezing and spray freeze-drying concepts, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
@@
Figure 2 summarizes the modeling strategies for the three different steps in lyophilization, including freezing, primary drying, and secondary drying. Published models mostly focus on primary drying because it is recognized as the most time-consuming, hazardous, and expensive step, and thus the first target for optimization and improvement. A variety of modeling strategies are available, ranging from the simplest lumped capacity model to the high-fidelity multidimensional model. Models for secondary drying have also been well developed as the final product quality (i.e., moisture content) is governed by this step. The mechanistic modeling of freezing receives the least attention in the lyophilization literature due to its complicated behavior, including subcooling and stochastic ice nucleation. As a result, the number of available models for freezing is more limited compared to that of the drying models.
@@
Modeling the freezing step is most complicated among the three steps of localization; the blue column in Fig. 2 highlights modeling strategies for freezing. Due to the limited number of freezing models, we expand our literature search beyond lyophilization to get more complete insights. A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [31]. Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails subcooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point. A more common but complicated strategy is to consider the subcooling period and incorporate the heat transfer associated with ice nucleation as done in Ref. [8]. Recently, Ref. [11] introduced a state-of-the-art freezing model which incorporates the stochastic nature of ice nucleation [32] and considers heat transfer among multiple vials in batch lyophilization. The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [12]. The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the mass transfer resistance during sublimation in primary drying. While some recent studies have built mechanistic models with the goal of predicting pore size distribution [10, 32], how to best model the specific complex interacting phenomena remains an open research question.
@@
Many models for primary drying are available in the literature (see the green column in Fig. 2). The simplest strategy is to use a lumped capacity model [22], which does not capture the effects of spatial gradients in the system. In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and the temperature gradient in the vertical direction is substantial. A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [33, 34, 8, 35]. By omitting mass transfer, the resulting model is simple and has only a few parameters, allowing the analytical solutions to be derived [35], and so the model can be implemented easily. However, as sublimation is a simultaneous heat and mass transfer process, the aforementioned models become inaccurate when the process is controlled by mass transfer. A clever way of incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of Ref. [18]). Of course, there are a number of models considering detailed heat and mass transfer, both in 1D [14, 15] and in higher dimensions [16]. Nevertheless, implementing those high-fidelity model for practical applications, e.g., state estimation and control, is not easy due to its high complexity, computational time, and number of parameters [18, 21]. Besides, the accuracy of complex models is not significantly different from 1D or simplified models in most cases [21]. Consequently, these high-fidelity models are not commonly used. Note that it is widely acknowledged that the amount of liquid removed via sublimation during primary drying is much higher than that of desorption, and so the effect of desorption can be omitted [15, 16, 18, 21].
@@
Secondary drying is similar to primary drying, with the liquid mainly removed via desorption instead of sublimation. Several modeling strategies are available, including a lumped capacity model [1], 1D heat transfer [25], 1D heat and mass transfer [15], and multidimensional models [16]. In the study by Ref. [25], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors. This observation is similar to the primary drying case, where a vertical thermal gradient is usually dominant.
@@
Here we select modeling strategies for our continuous lyophilization model that can capture critical phenomena in the process and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum. From the above discussion, the selected strategies are (1) freezing with nucleation and subcooling for the freezing step, (2) 1D heat transfer with mass transfer incorporated as a boundary condition for the primary drying step, and (3) 1D heat transfer with bound water desorption for the secondary drying step, which are also highlighted in Fig. 2. The models for continuous lyophilization developed in the subsequent sections are based on these strategies, with proper modifications for the suspended-vial configuration. Sections 3.2, 3.3, and 3.4 describe the models for freezing, primary drying, and secondary drying, respectively. Additionally, to supplement the mechanistic understanding of the proposed models, Section 3.5 specifically discusses modeling strategies and theories underpinning convection and thermal radiation, the key heat transfer mechanisms in suspended vials.
@@
Figure 3: Schematic diagram showing the mechanistic modeling of suspended vials for (A) freezing, (B) primary drying, and (C) secondary drying.
@@
For the design proposed by Ref. [6], suspended vials are cooled by using the cryogenic gas at a controlled temperature and flow rate. Besides, a dedicated chamber is added to control the nucleation temperature using the vacuum-induced surface freezing (VISF) technology [36]. With this setup, all surfaces of a vial experience a similar heat transfer condition, which is different from batch lyophilization where only the bottom surface is cooled by the cooling shelf. Given that the vial size is not large and heat transfer at all surfaces is almost identical, it is reasonable to use a lumped capacity model for the freezing step [11].
Our model (Fig. 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification, and (5) final cooling.
3.2.1 Preconditioning
In our model, the preconditioning step starts at $t_{0}$ and completes at $t_{\textrm{f1}}$. Preconditioning entails cooling the product in a vial such that the product temperature is uniform at the target value below the freezing point (subcooling). In terms of modeling, this is the simplest step because there is no phase change and nucleation, and so only sensible heat is important.
@@ 3.2.1 Preconditioning
where $T_{0}$ is the initial product temperature. During the freezing step in general, the amount of water changes with time due to phase transition, and so we denote $m_{\textrm{w0}}$ as the initial mass of water at $t_{0}$. On the other hand, the amount of solid does not change, meaning that $m_{\textrm{s}}$ is a constant. However, during the preconditioning step, there is no phase transition at all, so
@@
In some cases, experimental data are reported as the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$, where the subscript ‘l’ denotes the liquid phase (solvent + solute). In such cases, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated from $V_{\textrm{l}}$ and the densities $\rho_{\textrm{s}}$, $\rho_{\textrm{w}}$ (see Appendix A for the relevant equations).
The next step is to define the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$, which could vary among systems. For example, based on the data reported in Ref. [6], the bottom shelf is used to cool the gas inside the chamber, which subsequently cools the vials, hence natural convection. Besides, there exists thermal radiation between the vials and chamber walls. Therefore, both natural convection and radiation effects are modeled in this case. Consequently, the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are
$Q_{\textrm{s1}}=h_{\textrm{s1}}A_{z}(T_{\textrm{c}}-T),$ (4)
$Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T),$ (5)
$Q_{\textrm{s3}}=h_{\textrm{s3}}A_{z}(T_{\textrm{g}}-T)+\sigma A_{r}\mathcal{F}(T_{\textrm{c}}^{4}-T^{4}),$ (6)

### v2

# Mechanistic Modeling and Analysis of Transport Phenomena for Continuous Lyophilization

@@ Abstract
Lyophilization (aka freeze drying) is a typical process in pharmaceutical manufacturing used for improving the stability of various drug products, including its recent applications to mRNA vaccines. While extensive efforts have been dedicated to shifting the pharmaceutical industry toward continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in a batch mode. This article presents the first mechanistic model for a complete continuous lyophilization process, which comprehensively incorporates and describes key transport phenomena in all three steps of lyophilization, namely freezing, primary drying, and secondary drying. The proposed model considers the state-of-the-art lyophilization technology, in which vials are suspended and move continuously through the process. The validated model can accurately predict the evolution of critical process parameters, including the product temperature, ice/water fraction, sublimation front position, and concentration of bound water, for the entire process. Several applications related to model-based process design and optimization of continuous lyophilization are also demonstrated. The final model is made available in MATLAB and Julia as an open-source software package called ContLyo, which can ultimately be leveraged for guiding the design and development of future continuous lyophilization processes.
@@
Keywords— Lyophilization, Freeze Drying, Suspended Vials, Continuous Manufacturing, Pharmaceuticals
@@ Introduction
Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products in pharmaceutical manufacturing [1]. Various promising applications of lyophilization have been proven, including its recent applications that provide long-term stability for mRNA vaccines (e.g., for COVID-19) to allow for higher storage temperature [2, 3]. Advancement in lyophilization technology could therefore play a crucial role in the future of the pharmaceutical industry. A typical lyophilization process consists of three main steps, namely (1) freezing, (2) primary drying, and (3) secondary drying. During freezing, the product and liquid solvent (usually water) in a vial are cooled such that most of the liquid (free water) is frozen, with the remaining fraction (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals [4]. In primary drying, the free water in the form of ice crystals is removed via sublimation under vacuum. Finally, in secondary drying, the product is heated further such that the bound water can be removed via desorption.
@@
Although the current trends in the pharmaceutical industry largely focus on the adoption of continuous manufacturing, the majority of production-scale lyophilization is still being operated in a batch mode [5, 6], with most research efforts dedicated to process optimization, monitoring, and control to ensure that the final product quality meets the regulations [7, 6]. Conventional lyophilization of unit doses usually entails a batch of vials containing drug products situated on the cooling/heating shelf where the entire lyophilization process occurs. A number of continuous lyophilization concepts for pharmaceutical products have been proposed, including spray freeze drying [8] and spin freezing [5] (see a comprehensive review of continuous lyophilization technologies in Ref. [6]). A novel lyophilization technology has been recently proposed by Ref. [7], in which vials are suspended and continuously move along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
@@
Mathematical models have been widely used to assist the design, optimization, and control of lyophilization processes. Since the key phenomena in lyophilization are underpinned by heat and mass transfer theories, process models for lyophilization are mostly physics-based rather than data-driven. Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [9, 10, 11, 12, 13], primary drying [14, 15, 16, 17, 18, 9, 19, 20, 21, 4, 22], and secondary drying [14, 15, 16, 17, 18, 23, 4, 24, 25]. Only a few models for continuous lyophilization are available. For example, a model for the primary drying step in continuous drying of spin frozen vials was proposed by Ref. [26]. A model for the freezing step was also developed for spin freezing [27]. For spray freeze drying, Ref. [28] developed a detailed model that can predict the temperature of droplets during the cooling and freezing phases. Nevertheless, there is no model for the suspended-vial configuration. Besides, none of the published models for continuous lyophilization considers a complete lyophilization cycle (all three steps), which is critical for optimization and control of the entire continuous operation.
@@
This article presents the first mechanistic model for continuous lyophilization of suspended vials. The model is developed to capture important transport phenomena in the process, including cooling and stochastic/controlled ice nucleation during the freezing step, sublimation of ice crystals during the primary drying step, and desorption of bound water during the secondary drying step. The proposed model is validated and employed to predict the evolution of critical process parameters for the entire lyophilization process, which is subsequently demonstrated for model-based design and optimization.
@@
This article is organized as follows. Section 2 gives an overview of conventional (batch) and continuous lyophilization technologies. Section 3 extensively describes the development of our mechanistic model and its implementation, supported by a comprehensive review on a variety of modeling strategies used in the literature. Section 4 details the model validation and showcases various applications of the validated model. Finally, Section 5 summarizes the study and briefly discusses possible future work.
@@
The majority of lyophilization processes in the pharmaceutical industry have been operated in a batch mode for many decades. The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Fig. 1A). The bottom shelf, whose temperature can be manipulated, is used to cool the vials during the freezing step and heat the vials during the drying steps. In this case, the freezing and drying processes take place in the same space but at different times.
Figure 1: (A) Conventional batch lyophilization of unit doses. A number of vials are placed on the cooling/heating shelf. (B) Continuous lyophilization of suspended vials. A number of vials are suspended and move continuously through the lyophilizer.
@@
A comprehensive review of continuous lyophilization technologies is given in Ref. [6]. To summarize, the current continuous lyophilization technologies can be divided into four main categories. The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [29] or a revolving plate with holes [30]. Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with a high surface-to-volume ratio, greatly improving heat and mass transfer in the process [8, 28]. Both aforementioned approaches have been proposed for lyophilization of bulk material. For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [5].
Finally, the current state-of-the-art continuous lyophilization technology for unit doses has been recently proposed by Ref. [7]. This technology employs the suspended-vial configuration as illustrated in Fig. 1B, where all vials are suspended and continuously move through the freezing and drying chambers, without any contact between the vials and shelf. Unlike conventional lyophilizers, the freezing and drying processes occur in different locations/chambers of the equipment. This suspended-vial technology has several benefits over the existing ones. First, heat transfer in the equipment is uniform because every vial experiences the exact same heat transfer conditions along the way. In addition, there is a dedicated chamber for controlling the ice nucleation process to help reduce variation in the crystal structure caused by the stochastic nature of ice nucleation. Second, the process does not involve any complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control. Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in pharmaceutical applications. Finally, with the automatic filling and load-lock systems, the process is fully continuous.
The mechanistic model developed in this work is based on this suspended-vial configuration as illustrated in Fig. 1B, the current state-of-the-art continuous lyophilization technology for unit doses. Detailed modeling strategies and model equations are discussed in the next section (Section 3), while information and discussion related to the equipment design can be found in Ref. [7].
@@
This section details the development of a mechanistic model for continuous lyophilization of suspended vials (Fig. 1B). The section starts by discussing existing models and modeling strategies in the literature, and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
Review of the existing models and modeling strategies
Mechanistic modeling for conventional/batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry. Despite the fact that the models are for batch processes, modeling strategies underpinning those models can be used as a basis for the development of a continuous lyophilization model for suspended vials. While some past reviews briefly discuss the literature on mechanistic modeling of lyophilization [19, 4], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation. Hence, this section first provides a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes with the optimal modeling strategies used in this work. The published models for continuous lyophilization mentioned in Section 1 are based on the spin freezing and spray freeze-drying concepts, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
@@
Figure 2 summarizes various modeling strategies for the key steps of lyophilization, namely freezing, primary drying, and secondary drying. Published models mostly focus on primary drying because this step is recognized as the most time-consuming and expensive step, thus the first target for process optimization and improvement. A variety of modeling strategies are available, ranging from the simplest lumped capacity model to the high-fidelity multidimensional model. Models for secondary drying have also been well developed because the final product quality (i.e., moisture content) is governed by this step. The mechanistic modeling of freezing receives less attention in the lyophilization literature due to its complicated behavior, including supercooling and stochastic ice nucleation. As a result, the number of available models for freezing is more limited compared to that of the drying models.
Modeling strategies for the freezing step
@@
Modeling the freezing step is most complicated among the three steps of lyophilization; the blue column in Fig. 2 highlights notable modeling strategies for this step. Due to the limited number of freezing models, we expand our literature search beyond lyophilization to get more complete insights. A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [31]. Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails supercooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point [10, 12].
A more common but more complicated strategy is to consider the supercooling period and incorporate the heat transfer associated with ice nucleation as done in Ref. [9]. Recently, Ref. [12] introduced a state-of-the-art freezing model that incorporates the stochastic nature of ice nucleation [32] and considers heat transfer among multiple vials in batch lyophilization using a lumped capacity model. The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [13]. In Appendix A, we provide a detailed analysis using the relevant dimensionless group to compare between the lumped capacity model and model with thermal gradients and show that the lumped capacity model is sufficiently accurate for lyophilization of unit doses in general.
The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the solid structure and mass transfer resistance during sublimation in primary drying. While some recent studies have built mechanistic models with the goal of predicting pore size distribution [11, 32], how to best model the specific complex interacting phenomena remains an open research question.
Modeling strategies for the primary drying step
@@
Many models for primary drying are available in the literature (see the green column in Fig. 2). The simplest strategy is to use a lumped capacity model [22], which does not capture the effects of spatial gradients in the system. In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and hence the temperature gradient in the vertical direction is significant. A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [33, 34, 9, 35]. By omitting mass transfer, the resulting model is simple and has only a few parameters, allowing the analytical solutions to be derived [35], and so the model can be implemented easily. However, as sublimation is a simultaneous heat and mass transfer process, the aforementioned models become inaccurate when the process is controlled by mass transfer.
An approach to incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front/interface between the frozen and dried regions, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of Ref. [19]). In this case, the driving force for mass transfer is dependent on the saturation (equilibrium) pressure at the sublimation interface, which is a function of temperature. Hence, the model should be able to accurately predict the spatial variation of the product temperature, including the interface temperature, and hence modeling the heat transfer in the vertical direction is needed. Moreover, one of the key considerations during the drying steps is to ensure that the product temperature does not exceed the upper limit (e.g., collapse temperature and glass transition temperature) [1], and thus spatially distributed temperature data are valuable for accurately determining the maximum temperature in the product.
There are models that consider detailed heat and mass transfer by including full energy and continuity equations, both in 1D [15, 16] and in higher dimensions [17]. Implementing such high-fidelity models for practical applications, e.g., state estimation and control, is not easy due to their high complexity, computation time, and number of parameters [19, 4]. Besides, the accuracy of the complex models for primary drying is not significantly different from 1D or simplified models in most cases [4]. Consequently, these high-fidelity models are not commonly used.
Note that, in primary drying, the water vapor is removed via both sublimation (frozen region) and desorption (dried region). Nevertheless, it is widely acknowledged that the amount of water vapor removed via sublimation is much higher than that of desorption, and so the effect of desorption can be omitted [16, 17, 19, 4].
Modeling strategies for the secondary drying step
@@
Secondary drying is similar to primary drying, with the liquid mainly removed via desorption instead of sublimation. Several modeling strategies are available, ranging from the simplest lumped capacity model [1] to the high-fidelity model considering multidimensional heat and mass transfer in detail [17]. In Ref. [25], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors. Therefore, 1D modeling is preferable. Nevertheless, typical measurement techniques such as Karl Fischer titration measure the total amount of bound water, not the spatial variation. Consequently, a lumped capacity model is sometimes considered acceptable.
In primary drying, the drying rate is mainly governed by the rate of sublimation, which makes the modeling typically more straightforward than for secondary drying. In secondary drying, the two important mechanisms are (1) vapor phase transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product. Some previous models for secondary drying consider both phenomena [14, 15, 16, 17], whereas the more recently published models tend to consider only the desorption part [23, 4, 24, 25, 36]. Simulation results in these studies show that the model prediction is highly accurate even when the vapor phase transport is omitted, suggesting that desorption is the actual limiting step. In Ref. [37], a detailed experimental study was conducted, concluding that desorption is the rate-limiting step for mass transfer in secondary drying. Evidence from both simulation and experiment suggest that only the desorption part is necessary.
Although it is evident from both simulation and experiment in the literature that desorption is the limiting step, there is no systematic analysis from the theoretical perspective of transport phenomena. In Appendix A, we provide a simple and systematic way of analyzing the effect/contribution of each transport process via scale analysis, which can be used to justify the contribution of each transport process for different systems/conditions other than those considered in this work and in the literature.
Optimal modeling strategies
High-quality mechanistic modeling should balance between the model accuracy and complexity such that the resulting model can provide reliable results and be practically implemented for different purposes with ease, e.g., optimization, state estimation, and model-based control. High-fidelity models should have the best accuracy in theory, but these models typically entail a large number of equations and parameters. Simulating these model equations in real time could be challenging. Besides, incomplete knowledge of those parameters and associated uncertainty could negatively impact the model accuracy instead of improving it. On the other hand, low-fidelity models are much simpler to implement, but the model prediction might not be sufficiently accurate.
@@
Our goal is to select modeling strategies that can capture critical phenomena in the process and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum. From the discussion in Sections 3.1.13.1.3 and Appendix A, the selected modeling strategies are highlighted in Fig. 2. The freezing model relies on the lumped capacity method that considers nucleation and supercooling. The primary drying model simulates heat transfer in 1D (vertical direction) with mass transfer incorporated as a boundary condition at the sublimation front. Finally, the secondary drying model captures heat transfer in 1D (vertical direction) and desorption of bound water. The models for continuous lyophilization developed in the subsequent sections are based on these strategies, with proper modifications for the suspended-vial configuration. Sections 3.2, 3.3, and 3.4 describe the models for freezing, primary drying, and secondary drying, respectively. Additionally, to supplement the mechanistic understanding of the proposed models, Section 3.5 specifically discusses modeling strategies and theories underpinning convection and thermal radiation, the key heat transfer mechanisms in suspended vials.
@@
Figure 3: Schematic diagram showing the mechanistic modeling of continuous lyophilization via suspended vials for (A) freezing, (B) primary drying, and (C) secondary drying.
@@
For the design proposed by Ref. [7], suspended vials are cooled by using the cryogenic gas at a controlled temperature and flow rate. Besides, a dedicated chamber is added to control the nucleation temperature using the vacuum-induced surface freezing (VISF) technology [38]. With this setup, all surfaces of a vial experience a similar heat transfer condition, which is different from batch lyophilization where only the bottom surface is cooled by the cooling shelf.
Our model (Fig. 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification or ice formation, and (5) final cooling.
Preconditioning
In this model, the preconditioning step starts at $t_{0}$ and completes at $t_{\textrm{f1}}$. Preconditioning entails cooling the product in a vial such that the product temperature is uniform at the target value below the freezing point (supercooling). In terms of modeling, this step is the simplest because there is no phase change and nucleation, and so only sensible heat is important.
@@
where $T_{0}$ is the initial product temperature. During the freezing step in general, the amount of water changes with time due to phase transition, and so we denote $m_{\textrm{w,0}}$ as the initial mass of water at $t_{0}$. On the other hand, the amount of solid does not change, meaning that $m_{\textrm{s}}$ is a constant. However, during the preconditioning step, there is no phase transition at all, so
@@
In some cases, experimental data are reported as the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$, where the subscript ‘l’ denotes the liquid phase (solvent + solute). In such cases, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated from $V_{\textrm{l}}$ and the densities $\rho_{\textrm{s}}$, $\rho_{\textrm{w}}$ (see Appendix B for the relevant equations).
The next step is to define the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$, which could vary among systems. Here, we start by writing general expressions and then simplify them to match our system of interest. In suspended vials, there exists thermal radiation between the vials, chamber walls, and heating/cooling shelves. Therefore,
$Q_{\textrm{s1}}=h_{\textrm{s1}}A_{z}(T_{\textrm{g}}-T)+\sigma A_{z}\mathcal{F}_{\textrm{s1}}(T_{\textrm{u}}^{4}-T^{4}),$ (4)
$Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T)+\sigma A_{z}\mathcal{F}_{\textrm{s2}}(T_{\textrm{b}}^{4}-T^{4}),$ (5)
$Q_{\textrm{s3}}=h_{\textrm{s3}}A_{r}(T_{\textrm{g}}-T)+\sigma A_{r}\mathcal{F}_{\textrm{s3}}(T_{\textrm{c}}^{4}-T^{4}),$ (6)